# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v3

@@ 10 Compliance to AI Governance and Policy in Agentic AI Systems
As Agentic AI systems grow more autonomous and capable, ensuring they remain compliant with regulations and governed by robust policies becomes imperative. Unlike traditional AI, these multi-agent or highly autonomous systems operate with persistent memory and complex decision-making loops, which introduces new risks around safety, ethics, and legality. Recent industry surveys indicate that enterprise leaders rank security, trust, compliance, and oversight as top concerns when deploying agentic AI at scale. To address these concerns, governance of agentic AI spans three critical dimensions: adherence to regulatory standards, system-level auditability, and enforceable policy frameworks. Below, we expand on each dimension and highlight emerging best practices for secure and compliant agentic AI deployments.
@@ 10.1 Regulatory standards
Regulatory frameworks provide baseline requirements that all AI systems – including agent-based architectures – must meet. Key frameworks outline principles of transparency, accountability, safety, and fairness which are especially pertinent for autonomous multi-agent systems operating in high-stakes domains (e.g. healthcare, finance, transportation). Notable examples include:
  • NIST AI Risk Management Framework (AI RMF) [188], a voluntary U.S. framework emphasizing risk identification, mitigation, and trustworthy AI practices across the AI lifecycle. It calls for governance functions that ensure AI systems are transparent, accountable, and monitored for risks. Adopting the NIST AI RMF “Govern” and “Explain” functions helps organizations embed policies and documentation that align with trustworthiness goals.
  • EU AI Act [106], a comprehensive EU regulation (set to fully apply in 2026) that classifies AI systems by risk level. High-risk AI systems (a category that would likely include certain autonomous or agentic AI in critical fields) are subject to strict requirements: rigorous risk management, human oversight, transparency, and continuous monitoring. For example, the Act explicitly mandates ongoing risk assessment and traceability for high-risk AI, and it requires that such systems be designed with human-in-the-loop or human supervision capabilities (per Article 14 on oversight). Agentic AI developers must therefore build in features for traceability and human override to remain compliant with these upcoming rules.
  • Domain-Specific Regulations In addition to AI-focused standards, Agentic AI systems must comply with sectoral data and safety laws. For instance, an autonomous multi-agent healthcare assistant must adhere to health privacy laws like HIPAA [104], and any system handling personal data of EU citizens must follow GDPR [103] requirements for data protection by design. These regulations often require data minimization and user consent, which Agentic AI systems need to respect (e.g. by expiring or anonymizing stored user data to avoid retention of sensitive information).
@@ 10.2 Auditability
Auditability is critical for transparency and post-hoc accountability in agentic AI. Given the complexity of multi-agent decision-making (with potentially emergent or opaque interactions), organizations must be able to reconstruct and scrutinize the chain of events leading to any outcomes. Achieving robust auditability involves implementing the following practices:
  • Comprehensive Logging: Each decision, action, plan, and inter-agent message should be logged with rich metadata – including timestamps, the initiating agent’s identity/role, the context or prompt, and a rationale or justification for the action. Such detailed logs create a decision provenance trail, enabling auditors to trace how and why a particular outcome was reached.
  • Action Traceability: Techniques like decision provenance link each agent’s actions to preceding events or inputs, so one can follow the sequence of steps across agents. In complex agent workflows (e.g. a chain of a Planner agent delegating to a Coder agent, then to a Tester agent), traceability means every hand-off and decision node is recorded and can be reviewed in context. This helps identify where an error or policy violation might have occurred.
  • Role-Specific Audit Trails: Modern agentic frameworks often assign specialized roles to different agents (for example, in AutoGen or MetaGPT, agents may take on roles like Researcher, Coder, or Reviewer to collaborate on tasks). The audit system should capture actions at the granularity of each role. By tagging log entries with the agent role, one can detect anomalies or biases in specific functions – for instance, flagging if a “Reviewer” agent consistently approves outputs that the “Researcher” agent flagged as doubtful, or if a certain role’s actions lead to security issues
  • Immutable Logging: To ensure the integrity of audit logs (so that records cannot be tampered with after the fact), organizations are exploring append-only, tamper-proof logging mechanisms. For example, using blockchain or other immutable ledger technology to record multi-agent interactions can provide verifiable audit trails. This means any decision or message recorded in the system is cryptographically secured against alteration, which is valuable for compliance checks and forensic analysis in case of an incident.
Robust auditability not only helps with internal accountability but is also increasingly expected by regulators. Transparent record-keeping allows external auditors or regulators to verify compliance with regulations and ethical standards. In the event of a problematic outcome, detailed audit logs ensure that responsibility can be determined and corrective action taken. In fact, governance frameworks stress that organizations should implement such audit trails and explicable decision logs so that any harmful or biased result can be traced back and attributed appropriately. Overall, auditability in agentic AI builds the foundation for trust – both for internal stakeholders and for users or regulators – by making the system’s behavior observable and explainable after the fact.
@@ 10.3 Policy Enforcement
While regulations and audits provide oversight and after-the-fact analysis, policy enforcement mechanisms work in real-time to keep an Agentic AI system’s behavior within allowed bounds. Policy in this context refers to the codified rules and constraints that the AI agents must obey – covering operational limits, ethical guidelines, and security restrictions set by developers or governance boards. In practice, effective policy enforcement in agentic AI includes:
  • Orchestrator-Level Governance: Agentic AI systems typically have an orchestration layer or a meta-controller that coordinates multiple agents. This layer should be programmed to enforce global policies on what agents can do and under what conditions. For example, a policy might dictate that any plan an agent comes up with must receive approval from a supervisory agent (or a human) if it involves deleting or altering important data. The orchestration logic would then require that checkpoint before execution, overriding agent autonomy when necessary.
  • Memory and Data Retention Policies: To comply with privacy rules and avoid unchecked growth of an agent’s knowledge base, policies can enforce memory expiration or sanitization. Agents might be required to forget or redact certain information after a task is complete or after a time limit, preventing long-term retention of sensitive data. This guards against data misuse and helps comply with regulations like GDPR’s “data minimization” principle. For instance, if an agent system stores user queries or personal details, a policy could mandate purging those from the agent’s memory after fulfilling the query (unless explicitly allowed to retain).
  • Tool and Resource Access Control: Agentic AI often involves agents calling external tools, APIs, or databases to accomplish tasks. Role-Based Access Control (RBAC) and Attribute-Based Access Control (ABAC) are crucial for limiting what each agent (or agent role) is permitted to do. Under RBAC, each agent role has permissions aligned with its responsibilities – e.g. a “Coder” agent might be allowed to call a code compilation service but not to fetch customer data; a “Data Analyst” agent might read a database but not invoke external network calls. ABAC can add context-sensitive rules (attributes like time of day, the specific project, or the agent’s trust level) to further restrict actions.
  • Dynamic Sandboxing and Execution Constraints: Each agent’s actions (especially those that can change the environment or have side effects) should run in a sandbox or constrained context where possible. For example, if an agent is meant to write to a file, sandboxing can ensure it only writes to a designated directory and not to system-critical files. Similarly, agents might be prohibited from executing certain operations (like sending network requests to unapproved domains) through whitelists/blacklists coded into the environment. Dynamic sandboxing means these execution rules can adjust based on the agent’s role or current task, providing flexible yet controlled operation.
  • Real-time Monitoring and Intervention: Even with policies encoded, agents might encounter novel situations. Real-time monitoring systems serve as a safety net – they watch agent behavior and can halt or flag any activity that deviates from predefined ethical constraints or operational bounds. For instance, if an agent’s output is about to violate a known policy (say, it’s about to generate disallowed content or leak PII), a monitoring module can catch this and either censor the output or stop the agent’s process. Some approaches involve using formal verification and runtime checking tools (like TLA+ specifications or symbolic execution) to prove certain unsafe actions will not occur, or to continuously verify that invariants hold during execution.
10.4 Emerging Governance Practices
Because the field of Agentic AI is evolving rapidly, governance strategies are also adapting. Beyond the foundational controls above, organizations are adopting new best practices to strengthen oversight and ensure long-term compliance as these AI systems scale. Some of the emerging practices include:
  • AI Governance Boards: Many organizations are establishing dedicated governance boards or committees for AI oversight. These boards typically include stakeholders from legal, ethics, security, and domain experts. Their role is to continuously review the Agentic AI system’s objectives, assess risks, and update policies or strategies as needed. For example, a governance board might conduct regular audits of an AI’s decisions for bias or fairness, ensure alignment with the latest regulations, and approve any major changes to the system’s capabilities. This human layer of oversight creates accountability at the organizational level and sends a signal that the AI’s operations are being closely supervised by responsible parties.
  • Governance-as-Code: Inspired by the infrastructure-as-code paradigm in DevOps, governance-as-code means encoding governance policies and ethical rules directly into the software and configuration that drive the AI system. Instead of governance being only a manual or paperwork process, it becomes part of the system’s logic. For Agentic AI, this could involve using policy engines or rule-based systems that automatically check each agent action against a set of coded rules. For instance, a rule repository might specify “Agent of type X cannot use tool Y without permission Z” and this is enforced by the orchestration code.
  • Adaptive Governance Layers: Agentic AI systems may change over time – they might learn new skills, add more agents, or move into new deployment contexts. Adaptive governance refers to frameworks that evolve along with the AI system. This can include auto-updating rule sets or ML-driven policy adjustments. For example, if an agent starts operating in a new country, an adaptive governance layer might automatically incorporate that country’s AI regulations into the policy set. Or if a system update gives agents new capabilities, the governance rules adapt (or prompt human governors to adapt them) to cover those capabilities.
  • Standards Alignment and Certification: To have trust, organizations are aligning their Agentic AI governance with emerging international standards. A prime example is ISO/IEC 42001:2023 [189], and ISO/IEC 42005:2025 [101] the management system standard, which specify requirements for effective AI governance and risk management . ISO 42001 and 42005 emphasizes many of the same principles discussed above – such as transparency, accountability, bias mitigation, safety, and privacy in AI development and deployment. Forward-looking organizations are therefore proactively building those capabilities (e.g. implementing rigorous risk management pipelines, bias audits, etc.) so that by the time new laws fully kick in, their agentic AI will be in natural compliance.
By integrating these enforcement measures, Agentic AI systems can operate with much tighter alignment to intended guidelines and ethical norms. Policy enforcement translates high-level governance principles into day-to-day operational safeguards, reducing the chance of an autonomous agent straying outside its allowed domain or causing a compliance breach. Below, we summarize the core governance components required for secure and compliant Agentic AI deployment in Table 8.
@@ 11.5 Future Roadmap for Agentic AI TRiSM
Agentic AI Roadmap Scalable Agent Architectures Secure Multi-Agent Collaboration Explainable Multi-Agent Decisions Human-Centered Trust Design Autonomous Lifecycle Management Ethics & Governance Benchmarking & Evaluation Cognitive Capability Expansion
Figure 8: Agentic AI Roadmap: Mindmap Representation
@@
Policymakers and industry should collaborate to create regulatory sandboxes for multi-agent AI trials. These would be controlled environments where innovators can deploy Agentic AI systems under supervision, demonstrating TRiSM controls to regulators. Insights from such pilots can inform refinements in both technical standards and regulations.
There is much to learn from other high-stakes domains. For example, the safety engineering field (e.g. aerospace, automotive) has mature practices for redundant controls and failure mode analysis; these could inspire analogous practices in AI agent design. Likewise, ethics boards in biomedical research provide a template for AI ethics committees that review agent behaviors and approve high-risk deployments. We advocate establishing multidisciplinary governance boards that include ethicists, legal experts, domain specialists, and community representatives to oversee significant deployments of autonomous AI.
@@ 12 Conclusion
The rapid advancement of Agentic AI, driven by LLM-based multi-agent systems, marks a transformative shift in autonomous systems, enabling sophisticated collaboration, planning, and decision-making across high-stakes domains such as healthcare, finance, and public services. However, this evolution introduces complex challenges in trust, risk, and security management (TRiSM) that demand a tailored framework to ensure safe, transparent, and accountable operations. This review has explored these challenges, proposing a TRiSM framework structured around four pillars: Explainability, ModelOps, Security, Privacy and Governance, adapted for LLM-based AMAS. Key contributions include a novel risk taxonomy capturing unique threats like prompt injection, memory poisoning, and collusive failures, alongside two new metrics: the Component Synergy Score (CSS) and Tool Utilization Efficacy (TUE), which provide practical tools for assessing inter-agent collaboration and tool efficiency. The review also highlights state-of-the-art techniques, such as LIME, SHAP, and decision provenance, to enhance explainability, while advocating for robust security measures like prompt hygiene, sandboxing, and homomorphic encryption to mitigate vulnerabilities. Future work should prioritize advancing adversarial robustness, refining governance protocols, and developing standardized benchmarks for trustworthiness and coordination. By aligning Agentic AI development with TRiSM principles, researchers and practitioners can foster systems that are not only innovative but also responsible, ensuring their safe and ethical integration into society.

### v4

@@ 8 Compliance to AI Governance and Policy in Agentic AI Systems
Recent industry surveys indicate that enterprise leaders deploying Agentic AI at scale rank security, trust, compliance, and oversight among their top concerns. A survey [^14] TechRadar (2024) of IT professionals found that while 98% plan to expand AI agent usage, 96% view them as growing security threats, with only 54% having full visibility into their data access. In parallel, 67% of business leaders are increasing budgets for security oversight [^15] Cybersecurity Dive (2024), and 53% specifically cite data privacy and compliance as their primary challenge in scaling Agentic AI [^16] Cloudera (2024). Table LABEL:tab:agentic_governance operationalizes the governance controls into concrete implementations and evidence and next we detail the compliance to AI goverace and policy in Agentic AI systems.
@@ 8.1 Regulatory Standards
Regulatory frameworks provide baseline requirements that all AI systems, including agent-based architectures, must meet. Key frameworks include:
  • NIST AI Risk Management Framework (AI RMF) [171]. A voluntary U.S. framework organized around the Govern, Map, Measure, Manage functions. It emphasizes organizational governance, risk identification and measurement, and operational risk treatments. Mapping to the RMF also benefits from NISTIR 8312 (Explainable AI principles) for transparency obligations.
  • EU AI Act [125]. A comprehensive EU regulation with phased application: bans on certain practices apply after six months (Feb 2025), obligations for general-purpose AI begin Aug 2025, and most high-risk system obligations (e.g., risk management, documentation, human oversight, post-market monitoring) start Aug 2026. High-risk systems require ongoing risk assessment, technical documentation (Annex IV), logging/traceability (Arts. 11–12), data governance (Art. 10), human oversight (Art. 14), and robustness/accuracy (Art. 15).
  • ISO/IEC management and assessment standards. ISO/IEC 42001:2023 (AI management systems) sets organization-level governance requirements; ISO/IEC 42005:2025 (AI impact assessment) provides system-level impact-assessment guidance; ISO/IEC 23894 addresses AI risk management; ISO/IEC 24029-1 supports robustness evaluation; ISO/IEC TR 24027 addresses bias. These align well with AI RMF and the EU AI Act’s lifecycle duties.
  • Domain-specific laws. Sectoral/privacy regimes still apply: GDPR (e.g., data protection by design, Art. 25), HIPAA (45 CFR §164), and regional data-residency rules. Agentic systems should enforce data minimization, purpose limitation, and consent bases; consider DPIA for high-risk processing and AI impact assessments per ISO/IEC 42005 in parallel.
@@ 8.2 Auditability
Multi-agent systems can produce emergent and opaque behaviors; auditors must reconstruct the chain of decisions.
  • Comprehensive logging. Record prompts/contexts, plans, actions, tool calls, inputs/outputs, timestamps, and agent role/identity; capture rationales where feasible. Preserve hashes and version IDs for models, tools, and data snapshots.
[…]
  • Role-granular trails. Tag entries by agent role (e.g., Researcher, Coder, Reviewer) to localize anomalies and bias. Require reviewer countersignatures for high-impact actions.
  • Immutability. Use append-only, tamper-evident storage (e.g., WORM buckets, cryptographic sealing) for audit logs; rotate keys and attest integrity regularly.
@@ 8.3 Policy Enforcement
While regulations and audits provide oversight and after-the-fact analysis, policy enforcement mechanisms work in real-time to keep an Agentic AI system’s behavior within allowed bounds. Policy in this context refers to the codified rules and constraints that the AI agents must obey. In practice, effective policy enforcement in agentic AI includes:
  • Orchestrator-level controls. A meta-controller enforces allow/deny policies for plans and tool calls; escalate to human approval for sensitive operations (e.g., data deletion, external writes).
  • Memory & retention. Implement TTLs, redaction, and scoped memory (task- and tenant-scoped) to satisfy minimization and purpose limits; support subject-access and deletion.
  • Tool & data access. Enforce RBAC/ABAC for agents/roles; use per-tool API tokens, least privilege, and environment scoping (dev/stage/prod). Validate/normalize tool arguments.
  • Isolation. Run untrusted code/tooling in sandboxes (e.g., containers/VMs) with syscall/network/file egress policies; prefer hardware isolation (TEEs) for sensitive workloads.
  • I/O mediation. Apply content filters and egress controls on model outputs; mediate RAG inputs to prevent injection/exfiltration; detect sensitive data (PII/PHI, secrets) before release.
  • Real-time monitoring. Track anomaly signals (policy violations, unusual tool chains, data exfiltration patterns) and enforce runtime kill-switches, rollbacks, and incident triggers.
8.4 Documentation & Record-Keeping (Evidence)
To meet audit and regulator expectations, maintain:
  • Technical documentation (EU AI Act [1]): system purpose and risk class, training/eval data governance, performance, limitations, foreseeable misuse, human oversight measures.
  • Model/Data cards with lineage, licenses, and known hazards; evaluation reports with stress tests (robustness, red-teaming), calibration/fairness metrics, and residual risk.
  • Versioned artifacts: model versions, datasets, prompts, agents’ policies, and tool catalogs with change logs; signed releases and rollback points.
8.5 Post-Market Monitoring & Incident Response
Agentic systems must be monitored after deployment and respond to harm quickly.
  • Telemetry and drift. Continuously measure real-world performance, jailbreak/abuse rates, and data distribution shifts; trigger retraining or guard updates.
  • Serious-incident handling. Define thresholds, triage workflows, and reporting timelines; keep case files linking logs, affected users, and remediation.
  • Periodic reassessment. Re-run impact/risk assessments (ISO/IEC 42005 [129], ISO/IEC 23894 [130] ) when capabilities, data, or context change materially.
8.6 Supply-Chain & Secure Development (GenAI)
Strengthen provenance and development hygiene for agents and tools.
  • Secure SDLC for GenAI. Follow NIST SP 800-218 (SSDF) [^17] NIST SP 800-218 (SSDF) and the GenAI profile SP 800-218A  [^18] GenAI profile SP 800-218A for model/system development; integrate threat modeling for agent tool-chains and LLM-specific risks.
  • OWASP LLM risks. Mitigate prompt injection, insecure output handling, data poisoning, model DoS, and supply-chain risks using the OWASP [137] Top 10 for LLM applications.
  • Provenance. Track model/tool/dataset origins, licenses, and checksums; require attestations for third-party components and run dependency/secret scans in CI.
Bringing Agentic systems into compliance is not a single checklist; it is a lifecycle program that couples standards (NIST AI RMF, ISO/IEC 42001/42005), regulation (EU AI Act), secure development (NIST SP 800-218/218A), and runtime enforcement (policies, isolation, monitoring). This combination creates evidenceable trust and reduces the blast radius of inevitable failures.
@@ 9.5 Future Roadmap for Agentic AI TRiSM
Agentic AI Roadmap Scalable Agent Architectures Secure Multi-Agent Collaboration Explainable Multi-Agent Decisions Human-Centered Trust Design Autonomous Lifecycle Management Ethics & Governance Controllability Benchmarking & Evaluation Cognitive Capability Expansion
Figure 7: Agentic AI Roadmap: Mindmap Representation.
@@
Policymakers and industry should collaborate to create regulatory sandboxes for multi-agent AI trials. These would be controlled environments where innovators can deploy Agentic AI systems under supervision, demonstrating TRiSM controls to regulators. Insights from such pilots can inform refinements in both technical standards and regulations. There is much to learn from other high-stakes domains. For example, the safety engineering field (e.g. aerospace, automotive) has mature practices for redundant controls and failure mode analysis; these could inspire analogous practices in AI agent design. Likewise, ethics boards in biomedical research provide a template for AI ethics committees that review agent behaviors and approve high-risk deployments. We advocate establishing multidisciplinary governance boards that include ethicists, legal experts, domain specialists, and community representatives to oversee significant deployments of autonomous AI.
A cross-cutting concept relevant to both Governance and Lifecycle Management, controllability captures the ability of humans and oversight systems to steer or override agentic AI. Mechanisms include stop/pause functions, rollback gates role-based oversight, and auditability of inter-agent actions. In practice, controllability is critical to ensuring accountability and preventing runaway behaviors in multi-agent workflows.
@@ 10 Conclusion
The rapid advancement of Agentic AI, driven by AMAS, marks a transformative shift in autonomous systems, enabling sophisticated collaboration, planning, and decision-making across high-stakes domains such as healthcare, finance, and public services. However, this evolution introduces complex challenges in trust, risk, and security management (TRiSM) that demand a tailored framework to ensure safe, transparent, and accountable operations. This review has explored these challenges, proposing a TRiSM framework structured around five pillars: Explainability, ModelOps, Security, Privacy and their lifecycle Governance, adapted for LLM-based AMAS. Key contributions include a novel risk taxonomy capturing unique threats like prompt injection, memory poisoning, and collusive failures, alongside two new metrics: the Component Synergy Score (CSS) and Tool Utilization Efficacy (TUE), which provide practical tools for assessing inter-agent collaboration and tool efficiency. The review also highlights state-of-the-art techniques, such as LIME, SHAP, and decision provenance, to enhance explainability, while advocating for robust security measures like prompt hygiene, sandboxing, and homomorphic encryption to mitigate vulnerabilities. Future work should prioritize advancing adversarial robustness, refining governance protocols, and developing standardized benchmarks for trustworthiness and coordination. By aligning Agentic AI development with TRiSM principles, researchers and practitioners can foster systems that are not only innovative but also responsible, ensuring their safe and ethical integration into society.
Acknowledgements
Funding for the research was provided through Horizon Europe project AIXpert: An agentic, multi-layer, GenAI-powered backbone to make an AI system explainable, accountable, and transparent (ID:101214389).